At the local grocery store, apples are $1.20 per pound and oranges are $1.80 per pound. If the grocer wants to offer a 1.55-pound bag of mixed fruit for $1.62 per pound, how many pounds of apples must he use?

Found 2 solutions by josgarithmetic, richwmiller:
Answer by josgarithmetic(39618)

This is what you must understand:
http://www.algebra.com/my/mixture-price-two-part-both-parts-unknown.lesson?content_action=show_dev


The reason for the lesson is so you can solve more than just one specific exercise item ; and be able to solve EVERY problem that fits the same form.

Answer by richwmiller(17219)

a+b=1.55,
1.2*a+1.8*b=1.62*1.55
a=1.55-b
1.2*(1.55-b)+1.8*b=2.511
1.86-1.2b+1.8*b=2.511
0.6*b=0.651
b=1.085
a=1.55-b
a=0.465 lb. at $1.20
b=1.085 lb. at $1.80
check
1.2*0.465+1.8*1.085=1.62*1.55
0.558+1.953=2.511
2.511=2.511
ok